# Class-Aware PillarMix: Can Mixed Sample Data Augmentation Enhance 3D Object Detection with Radar Point Clouds?

### v1

@@ II Motivation
Figure 3: Examples of LiDAR-based MSDA methods applied to radar point clouds in Bosch Street dataset. The point clouds are aggregated from five sensors. In PolarMix, a 90-degree region along the azimuth axis is extracted and swapped between $s^{A}$ and $s^{B}$, while object-level points from $s^{B}$ are copied and rotated into the new sample $\tilde{s}$. In LaserMix, the point clouds are divided into six regions, and points from non-adjacent areas are combined from $s^{A}$ and $s^{B}$.
@@
    PolarMix [4] randomly divides the field of view into two regions based on the azimuth angle and then swaps the points within the selected angular interval. As shown in Fig. 3(a), a random 90-degree region is chosen for this exchange, with all points and corresponding labels being swapped. Additionally, points within bounding boxes in $s^{B}$ are copied and rotated into $\tilde{s}$ to enhance object diversity. Unlike LiDAR sensors, which offer uniform 360-degree detection, radar point clouds encounter irregularities in their angular distribution due to the radiation pattern, as shown in Fig. 2(a). This leads to significant variations in detection capability across the main lobe and side lobes, especially along the azimuth axis. Accordingly, substantial rotation operations such as PolarMix will disrupt the inherent angular distribution of the point cloud.
@@
As discussed above, the distinct point distribution patterns in radar point clouds, combined with their inherent sparsity and the use of multi-radar sensor fusion, differ significantly from 360-degree LiDAR data. Therefore, aforementioned methods developed for LiDAR data often fall short when applied to radar. Later results will reveal that some of these approaches not only underperform compared to general mixed-sample techniques like MixUp [11] but can even negatively impacting perception performance.
@@ III-C Class-aware $\beta$ distributions in $P_{c}$
1:Points and labels of two radar samples: $\{s^{A},\,y^{A}\},\{s^{B},\,y^{B}\}$; Pillar Size $\phi$; Pillar select ratio $\alpha$; MSDA probability $\delta$; MixUp $\beta$ distribution ${\color[rgb]{0.4660,0.6740,0.1880}\definecolor[named]{pgfstrokecolor}{rgb}{0.4660,0.6740,0.1880}\beta_{s}},\,{\color[rgb]{0.3010,0.7450,0.9330}\definecolor[named]{pgfstrokecolor}{rgb}{0.3010,0.7450,0.9330}\beta_{m}},\,{\color[rgb]{0.6350,0.0780,0.1840}\definecolor[named]{pgfstrokecolor}{rgb}{0.6350,0.0780,0.1840}\beta_{d}}$.
@@
8:         $\tilde{p}=MU(p_{A},\,p_{B},\,{\color[rgb]{0.4660,0.6740,0.1880}\definecolor[named]{pgfstrokecolor}{rgb}{0.4660,0.6740,0.1880}\beta_{s}})$, if $cls_{sparse}\;\textrm{in}\;p_{A}\,\textrm{or}\;p_{B}$
9:         $\tilde{p}=MU(p_{A},\,p_{B},\,{\color[rgb]{0.3010,0.7450,0.9330}\definecolor[named]{pgfstrokecolor}{rgb}{0.3010,0.7450,0.9330}\beta_{m}})$, elif $cls_{modarte}\;\textrm{in}\;p_{A}\,\textrm{or}\;p_{B}$
10:         $\tilde{p}=MU(p_{A},\,p_{B},\,{\color[rgb]{0.6350,0.0780,0.1840}\definecolor[named]{pgfstrokecolor}{rgb}{0.6350,0.0780,0.1840}\beta_{d}})$, elif $cls_{dense}\;\textrm{in}\;p_{A}\,\textrm{or}\;p_{B}$
11:         $\tilde{p}=MU(p_{A},\,p_{B},\,{\color[rgb]{0.3010,0.7450,0.9330}\definecolor[named]{pgfstrokecolor}{rgb}{0.3010,0.7450,0.9330}\beta_{m}})$, else
@@
16:    Update $\tilde{s},\,\tilde{y}$ by concatenating points from all $\tilde{p}$ and labels from both samples
@@ VI CONCLUSION
To the best of our knowledge, this work represents the first attempt to explore the application of MSDA to radar point clouds. We aim for this research to serve as a foundation and inspiration for future studies on MSDA in radar-based perception tasks. Moving forward, we plan to extend this approach to other contexts, such as semi-supervised learning and domain adaptation.

### v2

@@ II Related Works
Figure 3: Examples of LiDAR-based MSDA methods applied to radar point clouds in Bosch Street dataset. The point clouds are aggregated from five sensors. In PolarMix, a 90 region along the azimuth axis is extracted and swapped between $s^{A}$ and $s^{B}$, while object-level points from $s^{B}$ are copied and rotated into the new sample $\tilde{s}$. In LaserMix, the point clouds are divided into six regions, and points from non-adjacent areas are combined from $s^{A}$ and $s^{B}$.
@@
    PolarMix [4] randomly divides the field of view into two azimuth regions and swaps all points within the selected angular interval. As shown in Fig. 3(a), a random 90 region is chosen for this exchange, with all points and corresponding labels being swapped. Additionally, points within bounding boxes in $s^{B}$ are copied and rotated around the origin by a randomly selected angle within a predefined range into $\tilde{s}$ to enhance object diversity. Unlike LiDAR sensors, which offer uniform 360 detection, radar point clouds encounter irregularities in their angular distribution due to the radiation pattern, as shown in Fig. 2(a). This leads to significant variations in detection capability across the main lobe and side lobes, especially along the azimuth axis. Accordingly, substantial rotation operations such as PolarMix will disrupt the inherent angular distribution of the point cloud.
@@
As discussed above, the distinct point distribution patterns in radar point clouds, combined with their inherent sparsity and the use of multi-radar sensor fusion, differ significantly from 360 LiDAR data. Therefore, aforementioned methods developed for LiDAR data often fall short when applied to radar. Later results will reveal that some of these approaches not only underperform compared to general mixed-sample techniques like MixUp [11] but can even negatively impacting perception performance.
@@ III-C Class-aware $\beta$ distributions in $P_{c}$
1:Points and labels of two radar samples: $\{s^{A},\,y^{A}\},\{s^{B},\,y^{B}\}$; Pillar Size $\phi$; Pillar select ratio $\alpha$; MSDA probability $\delta$; MixUp $\beta$ distribution ${\color[rgb]{0.4660,0.6740,0.1880}\definecolor[named]{pgfstrokecolor}{rgb}{0.4660,0.6740,0.1880}\beta_{s}},\,{\color[rgb]{0.3010,0.7450,0.9330}\definecolor[named]{pgfstrokecolor}{rgb}{0.3010,0.7450,0.9330}\beta_{m}},\,{\color[rgb]{0.6350,0.0780,0.1840}\definecolor[named]{pgfstrokecolor}{rgb}{0.6350,0.0780,0.1840}\beta_{d}}$ for $cls_{moderate},cls_{sparse},cls_{dense}$.
@@
8:         if $cls_{sparse}\;\textrm{in}\;p_{A}\,\textrm{or}\;p_{B}$: $\tilde{p}=MU(p_{A},\,p_{B},\,{\color[rgb]{0.4660,0.6740,0.1880}\definecolor[named]{pgfstrokecolor}{rgb}{0.4660,0.6740,0.1880}\beta_{s}})$,
9:         elif $cls_{moderate}\;\textrm{in}\;p_{A}\,\textrm{or}\;p_{B}$: $\tilde{p}=MU(p_{A},\,p_{B},\,{\color[rgb]{0.3010,0.7450,0.9330}\definecolor[named]{pgfstrokecolor}{rgb}{0.3010,0.7450,0.9330}\beta_{m}})$,
10:         elif $cls_{dense}\;\textrm{in}\;p_{A}\,\textrm{or}\;p_{B}$: $\tilde{p}=MU(p_{A},\,p_{B},\,{\color[rgb]{0.6350,0.0780,0.1840}\definecolor[named]{pgfstrokecolor}{rgb}{0.6350,0.0780,0.1840}\beta_{d}})$,
11:         else: $\tilde{p}=MU(p_{A},\,p_{B},\,{\color[rgb]{0.3010,0.7450,0.9330}\definecolor[named]{pgfstrokecolor}{rgb}{0.3010,0.7450,0.9330}\beta_{m}})$
@@
16:    Update $\tilde{s},\,\tilde{y}$ by concatenating the points from every $\tilde{p}$ along with the labels of both samples
@@ VI CONCLUSION
To the best of our knowledge, this work represents the first attempt to explore the application of MSDA to radar point clouds. We aim for this research to serve as a foundation and inspiration for future studies on MSDA in radar-based perception tasks. Moving forward, we endeavor to explore strategies for optimizing the $\beta$ parameters during training. We also plan to extend this approach to other contexts, including semi-supervised learning and domain adaptation.